OpenStudy (anonymous):

An object A of mass 10 kg is moving with a velocity of 6 m/s. Calculate its kinetic energy and its momentum. If a constant opposing force of 20 N suddenly acts on A, find the times it takes to come to rest and the distance through which it moves

OpenStudy (anonymous):

\[E = \frac{ 1 }{ 2 } m v ^2 \] Is Kinetic Energy \[p = mv\] for momentum \[v = u + a t\] where \[u = 6\] \[a = -20\] \[v = 0\] Rearrange \[v = u + a t\] so that t is the subject